OpenStudy (anonymous):

There is a circle with a 16-inch chord. The midpoint of the chord is located 6 inches from the center of the circle. What is the length of the radius of the circle?

OpenStudy (anonymous):

Drawing this out you have a line going across the circle which has a midpoint which is 6 inches from the center of the circle. If you draw it you end up with a triangle which has the circle radius as a hypotenuse. The sides of the right triangle are 6 and 8 so the radius is: \[\sqrt{8^{2}+6^{2}}=\sqrt{64+36}=\sqrt{100}=10\]